OpenStudy (anonymous):

Elijah is braiding a cord that will be 28/24 of a yard long. he has finished 19/24 of a yard. How much more does he need to braid?

OpenStudy (whpalmer4):

Simple subtraction of fractions problem: \[\frac{28}{24} - \frac{19}{24} = \frac{28 - 19}{24} = \]

OpenStudy (anonymous):

ok, therefore the answer would be 4/24 right?

OpenStudy (whpalmer4):

Does 28 - 19 = 4?

OpenStudy (anonymous):

I mean 9/24

OpenStudy (whpalmer4):

Yes. Can you simplify that further?

OpenStudy (whpalmer4):

Depending on the circumstances, 9/24 might be the most convenient answer, but it's good to keep up the skill at reducing fractions, too :-)

OpenStudy (anonymous):

however they only have 1/24 or 23/24 or 1/6 or 1/12 to choose from.....

OpenStudy (whpalmer4):

Are you sure you copied the problem correctly? Are there units on the answers?

OpenStudy (anonymous):

These are the multipale chocie to choose from. 23/24yard 1/6 yard 1/12 yard 1/24 yard

OpenStudy (anonymous):

I came up with 9/24, but this is not one of my choices

OpenStudy (whpalmer4):

Okay, they are asking for an absolute amount, not a fraction of what's been done or anything like. None of those equal 3/8 of a yard, and that's exactly what 28/24 - 19/24 equals. Could that 8 actually be a 0 with some crap in the middle? That would give an answer of 1/24, which is among the choices.

OpenStudy (anonymous):

lol, no I copied the question from the source. that's what threw me. they want to know how much more does elijah need to braid. help, help

OpenStudy (whpalmer4):

Well, if the problem is correctly stated, the answer is 3/8 or 9/24 of a yard. If the answer has to be one of those choices, the problem is incorrectly stated. If you can't ask the person who wrote the problem, I would go with 1/24 in the hopes that it was a single-character typo in the problem.

OpenStudy (anonymous):

I am sorry it's 23/28

OpenStudy (anonymous):

I mean 23/24

OpenStudy (anonymous):

sorry, this is driving me crazy

OpenStudy (anonymous):

therefore, 4/24 needs further working right?

OpenStudy (texaschic101):

Is 23/24 the size of the braiding cord Elijah wants to make ?

OpenStudy (texaschic101):

4/24 = 1/6

OpenStudy (anonymous):

Elijah is braiding a cord that will be 23/24 of a yard long. he has finished 19/24 of a yard. How much more does he need to braid?

OpenStudy (anonymous):

how did you come up w/1/6

OpenStudy (texaschic101):

23/24 - 19/24 = 4/24 = 1/6 I just reduced 4/24 to get 1/6

OpenStudy (anonymous):

ok, how did you reduce, I am trying to teach my granddaughter how to reduce it........

OpenStudy (texaschic101):

what number goes into 4 and 24 evenly ? 4 does. 4 goes into 4 one time and 4 goes into 24, six times.

OpenStudy (anonymous):

You are so special for taking out this time with me, thank you.

OpenStudy (texaschic101):

do you understand how to reduce ? because if you don't, I can give you more examples

OpenStudy (anonymous):

one more example wouldn't hurt, thanks - teach me,

OpenStudy (texaschic101):

3/18 what goes into 3 and 18 evenly ?

OpenStudy (anonymous):

3 goes into 3 one time, and 3 goes into 18 six times. so would the answer be 3/6

OpenStudy (texaschic101):

no......it would be 1/6 not 3/6

OpenStudy (anonymous):

oh, thats right because 4 goes into 4 one time. Ok, give me one more.

OpenStudy (texaschic101):

6/18

OpenStudy (anonymous):

1/3

OpenStudy (whpalmer4):

Ah, good eyes, spotting that it is supposed to be 23/24!

OpenStudy (anonymous):

is 1/3 correct?

OpenStudy (texaschic101):

correct. It is easier to use the largest number that will go into both numbers, but even if it isn't the largest number, it will still work. For instance if you used 6...6/18 = 1/3 and if you used 3....6/18 = 2/6 = 1/3. By not using the largest number, this just means you might have to reduce more then once. Now if you get an answer like 5/7. What goes into 5 and 7 evenly ? Nothing does, so this fraction will not reduce. Understand ?

OpenStudy (whpalmer4):

To reduce a fraction, in general, rewrite it as all of its prime factors. Prime numbers are those such as 2, 3, 5, 7, 11, 13, etc. which don't have any divisors except 1 and themselves. 2 is the only even prime number. So, to reduce \(\dfrac{4}{24}\) we could write \[\frac{4}{24} = \frac{2*2}{2*2*2*3}\] and then cancel matching items from the top and bottom giving us \[\frac{1}{2*3} = \frac{1}{6}\]

OpenStudy (anonymous):

lets try this one 5/6 - 1/6 so this would be 4/6 = 1/6 right? or not?

OpenStudy (anonymous):

or 2/3

OpenStudy (whpalmer4):

\[\frac{5}{6}-\frac{1}{6} = \frac{4}{6} = \frac{2*2}{2*3} = \frac{2}{3}\]

OpenStudy (texaschic101):

2/3 is correct....2 goes into 4 two times and 2 goes into 6 three times

OpenStudy (anonymous):

now how about if its like this? 1/12 + 3/4 =

OpenStudy (texaschic101):

turn mixed numbers into improper fractions, then add. Do you know how to do that ?

OpenStudy (texaschic101):

oops read it wrong

OpenStudy (anonymous):

no, because I would have just said 4/16

OpenStudy (texaschic101):

1/12 + 3/4 To add fractions you have to have the same denominator. If the denominators are not the same, you have to make them the same. What is the lowest number that 4 and 12 will go into ? 12 is the lowest number. So, you have to make the denominator 12. So, in this case 1/12 stays the same but 3/4 will change. Look at 3/4 and you want 12 to be the denominator.....4 goes into 12 how many times ? 3 times. Now take that 3 and multiply it by the numerator (3) and put that answer over 12. So, now you have 9/12. Your problem now is 1/12 + 9/12 = 10/12 = 5/6

OpenStudy (anonymous):

wow, stick a fork in me becasue I am done. I have copied all your notes to study, so that I can break it down to him, again thank you so much, I need a cup of coffee after that. lol you are very special. again thanks

OpenStudy (texaschic101):

that is what we are here for....if you have any more problems, we are here to help :)

OpenStudy (texaschic101):

Remember this as well.....you have to have a common denominator in both adding and subtracting fractions, but not when multiplying or dividing fractions.

OpenStudy (whpalmer4):

You need to find a common denominator. Ideally, that common denominator will the Least Common Multiple (LCM) of all of the denominators in the problem. To find the LCM, you take the prime factors of each denominator: \[4 = 2*2 = 2^2\]\[12 = 2*2*3 = 2^2*3\] Then you multiply the highest power of each different factor: \[2^2*3 = 12\] This seems like more work than just spotting that 12 = 3*4, but when you have a big ugly number, it saves work. For example, \[\frac{1}{8}+\frac{1}{9}+\frac{1}{10}+\frac{1}{11}\] \[8 = 2^3\]\[9= 3^2\]\[10=2*5\]\[11=1*11\] Now we combine the highest power of each factor: \[2^3*3^2*5*11 = 8*9*5*11 = 3960\]So our answer will be \[\frac{x}{3960}\] (for the appropriate value of \(x\).

OpenStudy (whpalmer4):

As an exercise, try working out that fraction. You'll multiply the 1/8 by 9*10*11/9*10*11, the 1/9 by 8*10*11/8*10*11, and so on, then add the results.

OpenStudy (texaschic101):

good explaining whpalmer4 :)

OpenStudy (whpalmer4):

That looks like such an innocuous little fraction, doesn't it? :-)

OpenStudy (texaschic101):

lol....it sure does

OpenStudy (texaschic101):

Darn....I should have explained to her what a fraction like 4/4 equals. Do you think she will know that ?

OpenStudy (whpalmer4):

One more thing: If you are doing an algebra problem, not just adding a few fractions, you can make life much easier by not doing fractions at all! For example, say our fiendish fraction problem was \[\frac{x}{8}+\frac{1}{9}+\frac{1}{10}+\frac{1}{11} = 1\]We could find the common denominator and convert everything over to it, but that's a bunch of work. Instead, we just multiply both sides of the equation by all of the denominators: \[(8*9*10*11)(\frac{x}{8}+\frac{1}{9}+\frac{1}{10}+\frac{1}{11} )= 1(8*9*10*11)\]which simplifies to \[(9*10*11)x + 8*10*11 + 8*9*11 + 8*9*10 = 8*9*10*11\]\[990x + 2392 = 7920\]and it's just a short hop and a skip to \[x = \frac{2764}{495}\]

OpenStudy (whpalmer4):

This works because the product of all of the denominators is clearly a multiple of all of them. If the numbers are all relatively prime (they share no factors other than 1), this will in fact be the least common multiple (or least common denominator).